# No-prior Bayes reIMagined: probabilistic approximations of possibilistic inferential models

### v2

@@ 4.1 Agrees with Bayes in group invariant models
Let $\mathscr{G}$ denote a group of transformations from $\mathbb{X}$ to $\mathbb{X}$, with function composition $\circ$ as the binary operation. As is customary, I’ll write $gx$ for the image of $x\in\mathbb{X}$ under transformation $g\in\mathscr{G}$; and if $g_{1}$ and $g_{2}$ are two group elements, then $g_{1}\circ g_{2}$ denotes their composition. Since $\mathscr{G}$ is a group, it’s associative, i.e., $g_{1}\circ(g_{2}\circ g_{3})=(g_{1}\circ g_{2})\circ g_{3}$ for all $g_{1},g_{2},g_{3}\in\mathscr{G}$, it contains the identity transformation $e$, and for every $g\in\mathscr{G}$, there exists an inverse $g^{-1}\in\mathscr{G}$ such that $g\circ g^{-1}=g^{-1}\circ g=\text{identity}$. Examples of transformation groups include translation, scaling, rotation, permutation, etc.
@@
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. More details on group invariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix B.
@@
When the model is invariant in the sense described above, there are standard and broadly agreed upon no-prior Bayes solutions, namely, those based on combining the likelihood with the so-called right Haar prior using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others. In Martin, 2023a , I showed that, under the above conditions, there is a close connection between the possibilistic IM and this standard no-prior Bayes solution. In the present context, the natural interpetation of the result established there is that the standard no-prior Bayes solution—namely, the Bayes posterior with respect to the invariant right Haar prior—is an inner probabilistic approximation of the possibilistic IM. Further details, including a new proof, are provided in Appendix B.
Theorem 3 (Martin, 2023a ).
@@
For an invariant statistical model, under the conditions laid out in Appendix B, the Bayesian posterior distribution relative to the right Haar prior is an inner probabilistic approximation of the possibilistic IM.
@@
That is, under an invariant statistical model, the familiar Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. This justifies my previous claim that the Bayes posterior in the gamma illustration, as shown in Figure 2 above, is an inner probabilistic approximation. This connection also serves as a strong proof of concept for the no-prior Bayes solution that I’m proposing here: that is, the right Haar prior Bayesian solutions in such problems are reasonable and widely used, so what I’m proposing must be equally reasonable and usable. Finally, it also helps to explain certain phenomena from the Bayesian perspective, namely, why the Bayesian/fiducial solutions in these problems are exactly probability matching.
@@
A point made above that deserves to be mentioned again is that the inner probabilistic approximations are not unique. So, the Bayes solution here might not exactly correspond to the generically recommended strategy that defines $\mathsf{K}_{x}^{\beta}$ to be uniform on $\partial C_{\beta}(x)$; the illustration in Figure 2 is such a case. In fact, when a direct connection between a Bayesian solution and the IM’s inner probabilistic approximation can be made, then I might recommend the former since its computation is more familiar than the latter’s. But, like in Figure 2, the difference between these inner probabilistic approximations will be negligible and will vanish as the information in data increases; see Section 4.2.
@@
As an illustration, consider a case involving directional data, where the observations correspond to directions—points on a unit circle in the plane—or simply angles with respect to a reference direction. Real-world applications involving these kind of data include wind and animal movement studies; see Mardia and Jupp, (2000) for a comprehensive treatment. More generally, direction measurements can be represented by points on the surface of a hypersphere. This kind of data is common in astronomy, where the position of planets, stars, etc. can be described by a point on the celestial sphere. The present example involves $n=25$ such planar directional observations, plotted in Figure 3. These data are simulated from a model very similar ot that used to analyze a very small data set on roulette wheel spins from Example 1.1 in Mardia and Jupp, (2000). The posited model is the so-called von Mises distribution, parametrized by a mean direction/angle $\theta$ and a concentration parameter $\kappa$. Further details about this model and the corresponding IM/Bayes/fiducial solutions are given in Appendix B.2. Here I’ll take the concentration parameter $\kappa=4$ to be fixed and known. I’ll also interpret the data as angles (radians) so that the parameter $\theta$ acts like a location paramter. With this simplification, the $\theta$-only von Mises model is invariant with respect to addition modulo $2\pi$. Therefore, Theorem 3 implies that the Bayes/fiducial distribution is an inner probabilistic approximation of the possibilistic IM. Figure 4 shows, first, a plot of the possibilistic IM’s contour for these data, with its peak at the maximum likelihood estimator, $\hat{\theta}_{x}=1.35$ and, second, a histogram of samples drawn from the inner probabilistic approximation (with $\mathsf{K}_{x}^{\beta}$ a uniform distribution) and the no-prior Bayesian posterior density overlaid. The similarities provided a visual confirmation of the claim in Theorem 3.
@@
Figure 3: Plot of the directional data analyzed in Section 4.1. Points on the circle determine the direction. The small closed circle denotes the sample mean in Cartesian coordinates and the corresponding angle determined by that point, $\hat{\theta}_{x}=1.35$, is the maximum likelihood estimator of the mean angle. The length of that vector, 0.87, is a to-be-conditioned-on ancillary statistic; see Appendix B.2.
@@ 4.2 Asymptotic normality and efficiency
This section presents a version of the celebrated Bernstein–von Mises theorem for the possibilistic IM, which establishes its asymptotic normality and efficiency. That is, for large $n$, the possibilistic IM’s contour resembles a Gaussian possibility contour, defined below, with covariance matrix matches the familiar Cramér–Rao lower bound. This implies a traditional Bernstein–von Mises theorem for the proposed inner probabilistic approximation and, therefore, it agrees with any other reasonable no-prior Bayesian solution asymptotically. Moreover, this near-Gaussian form of the IM output offers some valuable insights and simplifications relevant to computing the inner probabilistic approximation; see Section 5. The first order of business is to define Gaussian possibility.
Let $g_{m,v}$ denote the $D$-dimensional Gaussian probability density function, parametrized by a mean vector $m\in\mathbb{R}^{D}$ and a covariance matrix $v\in\mathbb{R}_{+}^{D\times D}$. Define the corresponding Gaussian possibility measure as the outer possibilistic approximation (e.g., Dubois and Prade, 1990) of ${\sf N}_{D}(m,v)$, i.e., the possibility measure with the smallest credal set that contains ${\sf N}_{D}(m,v)$. Analogous to the formula (2) presented in Section 2, this is just the probability-to-possibility transform applied to the aforementioned Gaussian distribution, and the corresponding possibility contour is
$\gamma_{m,v}(y)=\mathsf{P}\{g_{m,v}(Y)\leq g_{m,v}(y)\},\quad y\in\mathbb{R}^{D},\quad Y\sim{\sf N}_{D}(m,v).$
This can easily be rewritten in a beneficial and effectively closed-form expression,
@@
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. Write $\overline{\Gamma}_{m,v}$ for the corresponding possibility measure obtained by maximizing the contour as described above. Other authors have used a different form of Gaussian possibility (e.g., Denœux, 2023b ; Denœux, 2023a ), but I find the one above appealing and convenient because of the following duality between ${\sf N}_{D}(m,v)$ and $\overline{\Gamma}_{m,v}$: the former is the inner probabilistic approximation of the latter, and the latter is the outer possibilistic approximation of the former.
For some intuition about how this is related to the possibilistic IMs, consider a simple Gaussian location problem, $X\sim\mathsf{P}_{\Theta}={\sf N}_{D}(\Theta,v)$, with known covariance matrix $v$. Then the relative likelihood is $R(x,\theta)=\exp\{-\frac{1}{2}(x-\theta)^{\top}v^{-1}(x-\theta)\}$, quadratic on the log scale, and, since $\|X-\theta\|^{2}\sim{\sf ChiSq}(D)$ under $\mathsf{P}_{\theta}$, it follows that possibilistic IM’s contour is
$\pi_{x}(\theta)=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}=1-F_{D}\bigl{(}(x-\theta)^{\top}v^{-1}(x-\theta)\bigr{)},\quad\theta\in\mathbb{R}^{D}.$
@@
$\pi_{x^{n}}(\theta)\approx 1-F_{D}\bigl{(}(\hat{\theta}_{x^{n}}-\theta)^{\top}J_{x^{n}}(\hat{\theta}_{x^{n}}-\theta)\bigr{)},\quad\theta\in\mathbb{R}^{D},$
where $\hat{\theta}_{x^{n}}$ is the maximum likelihood estimator and $J_{x^{n}}$ is the observed Fisher information matrix. The reader will notice that the right-hand side above is exactly the Gaussian possibility contour with mean vector $\hat{\theta}_{x^{n}}$ and covariance matrix $J_{x^{n}}^{-1}$, hence the connection between the possibilistic IM and the Gaussian possibility is very natural and holds quite generally, at least approximately. Making this connection rigorous, however, requires considerable care, both in formulating the proper regularity conditions and in suitably bounding all the error terms. Along these lines, under regularity conditions comparable to those advanced in Le Cam, (1956, 1960, 1970)—see, also, van der Vaart, (1998) and Bickel et al., (1998)Martin and Williams, (2025) proved the following novel Bernstein–von Mises theorem for the possibilistic IM.
@@
Under the regularity conditions presented in Appendix C, the possibilistic IM’s contour $\pi_{X^{n}}$ satisfies
$\sup_{\theta\in\mathcal{T}}\bigl{|}\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathcal{T}$ is a fixed but arbitrary compact subset of $\mathbb{T}$, $\gamma_{X^{n}}(\theta)=\gamma_{\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1}}(\theta)$ is the centered and scaled Gaussian possibility contour, $I_{\Theta}$ the Fisher information, and $\Delta_{\Theta}(X^{n})$ a sequence of random variables that converges in distribution to ${\sf N}_{D}(0,I_{\Theta}^{-1})$. If, in addition, the likelihood function is twice differentiable, then
$\sup_{u\in\mathcal{U}}\bigl{|}\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}u)-\gamma(u)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
@@
Note that the covariance matrix in the Gaussian limit agrees with the Cramér–Rao lower bound. It’s in this sense that the possibilistic IM is asymptotically efficient and, hence, there’s no cost in terms of efficiency for the IM’s exact validity (and imprecision).
A direct consequence of Theorem 4 is that the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the classical, Bayesian Bernstein–von Mises theorem. The proof is in Appendix C.
@@
Under the conditions of Theorem 4, the approximation $\mathsf{Q}_{X^{n}}^{\star}$ of the possibilistic IM satisfies a probabilistic Bernstein–von Mises theorem, i.e.,
$d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}}):=\sup_{H}\bigl{|}\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr{|}\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$},$
where $\mathsf{G}_{X^{n}}={\sf N}_{D}(\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1})$ is the approximating Gaussian distribution and the supremum is over all Borel measurable subsets $H$ of $\mathbb{T}$.
I’ll give a quick sketch of the proof for a scalar parameter $\Theta$. For a generic value $\theta$, if the level set $C_{\alpha}(x^{n})$ is the interval $[a_{\alpha}(x^{n}),b_{\alpha}(x^{n})]$, then it’s easy to verify that
$\{a_{\alpha}(x^{n})\geq\theta\}\;\text{or}\;\{b_{\alpha}(x^{n})\leq\theta\}\iff\pi_{x^{n}}(\theta)\leq\alpha.$
Using the mixture characterization of the inner probabilistic approximation, one can verify that the cumulative distribution function of $\Theta$, relative to $\mathsf{Q}_{x^{n}}^{\star}$, at $\theta$ is
[…]
=$=\begin{cases}\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta\leq\hat{\theta}_{x^{n}}$}\\ 1-\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta>\hat{\theta}_{x^{n}}$}.\end{cases}$
Switching to local coordinates, i.e., $\theta=\hat{\theta}_{x^{n}}+J_{x^{n}}^{-1/2}z$ for $z$ fixed (or bounded in $n$), the case $\theta\leq\hat{\theta}_{x^{n}}$ corresponds to $z\leq 0$ and Theorem 4 implies
@@
$=\tfrac{1}{2}\,\mathsf{P}(Z^{2}\geq z^{2}),\quad\text{for $Z\sim{\sf N}(0,1)$}$
@@
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, identical to the classical Bernstein–von Mises theorem.
A point that follows immediately from the above sketch is that assigning probability $\frac{1}{2}$ to the boundary points in the inner probabilistic approximation is crucial; of course, any choice that assigns probability to the boundary points that converges to $\frac{1}{2}$ works fine too. More generally, since $\mathsf{G}_{X^{n}}$ is an inner probabilistic approximation to the possibility measure with contour $\gamma_{X^{n}}$, and corresponds to taking a uniform kernel on the level sets of $\gamma_{X^{n}}$, it’s necessary that the choice of kernel defining $\mathsf{Q}_{X^{n}}^{\star}$ at least be asymptotically uniform. This implies that the inner probabilistic approximation’s non-uniqueness vanishes as $n\to\infty$. This observation adds further support for my recommendation above to adopt the uniform kernel in the inner probabilistic approximation.

### v3

@@ 4.1 Agrees with Bayes in group invariant models
Let $\mathscr{G}$ denote a group of transformations from $\mathbb{X}$ to $\mathbb{X}$, with function composition $\circ$ as the binary operation. As is customary, I’ll write $gx$ for the image of $x\in\mathbb{X}$ under transformation $g\in\mathscr{G}$. Since $\mathscr{G}$ is a group, it’s associative, i.e., $g_{1}\circ(g_{2}\circ g_{3})=(g_{1}\circ g_{2})\circ g_{3}$ for all $g_{1},g_{2},g_{3}\in\mathscr{G}$, it contains the identity transformation, and for every $g\in\mathscr{G}$, there exists an inverse $g^{-1}\in\mathscr{G}$ such that $g\circ g^{-1}=g^{-1}\circ g=\text{identity}$. Examples of transformations include translation, scaling, rotation, permutation, etc.
@@
For example, if the distribution of $X$ depends on a location parameter $\theta$, then the distribution of $X+\tau$ depends on parameter $\theta+\tau$. When the statistical model $\{\mathsf{P}_{\theta}:\theta\in\mathbb{T}\}$ satisfies (10), it’s called an invariant statistical model—that is, the model for $X$ is also the model for $gX$, with $g\in\mathscr{G}$. Let $\overline{\mathscr{G}}$ denote the collection of transformations $\bar{g}$ on $\mathbb{T}$, corresponding to the mappings $g\in\mathscr{G}$ and parameters $\theta\in\mathbb{T}$; it’s easy to check that $\overline{\mathscr{G}}$ is a group too. More details on equivariant models can be found in Fraser, (1968), Eaton, (1989), and Schervish, (1995, Ch. 6), among others. The necessary technical background is given in Appendix C of the supplementary material.
@@
For invariant models, a standard and broadly agreed upon no-prior Bayes solution $\mathsf{Q}_{x}^{\rho}$ is that based on combining the likelihood with a so-called right Haar prior $\rho$ using Bayes’s formula (e.g., Kass and Wasserman, 1996). This also agrees with the solution obtained via Fisher’s fiducial argument, among others; see Lindley, (1958).
To clarify the scope of the following result, I’m assuming that the group $\mathscr{G}$ with respect to which the model is invariant is unique. In that case, under very mild conditions, the right Haar measure $\rho$ and the corresponding Bayes solution $\mathsf{Q}_{x}^{\rho}$ is unique too. But there are models, including the multivariate Gaussian models studied in Sun and Berger, (2007), that are invariant with respect to more than one group. So, even though each group has a unique right Haar prior, the non-uniqueness of the group implies non-uniqueness of the right Haar prior and corresponding Bayes solution. This latter case is more nuanced and I’ll discuss these details in Appendix C.
Theorem 3.
@@
If the model is invariant with respect to a unique group $\mathscr{G}$, then, under the conditions laid out in Appendix C, the posterior $\mathsf{Q}_{x}^{\rho}$ relative to the right Haar prior $\rho$ is an inner probabilistic approximation of $\mathsf{\Pi}_{x}$.
@@
That is, for invariant models, the standard no-prior Bayes solutions are inner probabilistic approximations of the possibilistic IM. This justifies my previous claim that the Bayes posterior in the gamma illustration, in Section 3.3, is an inner probabilistic approximation. This also serves as a general proof of concept for the no-prior Bayes solution that I’m proposing here: that is, at least in cases where the group is unique, the corresponding Bayes solutions are reasonable and widely used, so what I’m proposing must be equally reasonable and usable.
@@
Remember that inner probabilistic approximations are not unique. So, the right Haar prior Bayes solutions might not exactly match my recommended strategy that defines $\mathsf{K}_{x}^{\beta}$ to be uniform on $\partial C_{\beta}(x)$; the illustration in Figure 2 is such a case. In fact, when there’s a Bayesian solution that’s an inner probabilistic approximation of the IM, then I might recommend the former since its computation is more familiar than the latter’s. But, like in Figure 2, the difference between each inner probabilistic approximations will be negligible and will vanish as the information in data increases; see Section 4.2.
@@
As an illustration, suppose that observations correspond to directions—points on a unit circle in the plane—or simply angles with respect to a reference direction. Real-world applications involving these kind of data include wind and animal movement studies; see Mardia and Jupp, (2000) for a comprehensive treatment. More generally, direction measurements can be represented by points on the surface of a hypersphere. This is common in astronomy, where the position of a planet or star can be described by a point on the celestial sphere. The present example involves $n=25$ such planar directional observations, plotted in Figure 3. These data are simulated from the so-called von Mises distribution, parametrized by a mean direction/angle $\theta$ and a concentration parameter $\kappa$. Further details about this model and the corresponding solutions are given in Appendix C.2 in the supplement. Here I’ll take the concentration parameter $\kappa=4$ to be fixed and known. I’ll also interpret the data as angles (radians) so that the parameter $\theta$ acts like a location paramter. With this simplification, the $\theta$-only von Mises model is invariant with respect to addition modulo $2\pi$. Therefore, Theorem 3 implies that the Bayes solution is an inner probabilistic approximation of the possibilistic IM. Figure 4 shows, first, a plot of the possibilistic IM’s contour for these data, with its peak at the maximum likelihood estimator, $\hat{\theta}_{x}=1.35$ and, second, a histogram of samples drawn from the inner probabilistic approximation (with $\mathsf{K}_{x}^{\beta}$ a uniform distribution) and the no-prior Bayesian posterior density overlaid. The visual similarities provide confirmation of the claim in Theorem 3.
@@
Figure 3: Plot of the directional data analyzed in Section 4.1. Points on the circle determine the direction. The small closed circle denotes the sample mean in Cartesian coordinates and the corresponding angle determined by that point, $\hat{\theta}_{x}=1.35$, is the maximum likelihood estimator of the mean angle. The length of that vector, 0.87, is a to-be-conditioned-on ancillary statistic; see Appendix C.2
@@ 4.2 Asymptotic normality and efficiency
Next is a version of the celebrated Bernstein–von Mises theorem for the possibilistic IM, establishing its asymptotic normality and efficiency. That is, for large $n$, the possibilistic IM’s contour resembles a Gaussian possibility contour, defined below, with covariance matrix that matches the Cramér–Rao lower bound. This implies a traditional Bernstein–von Mises theorem for the proposed inner probabilistic approximation and, therefore, it agrees with any other reasonable no-prior Bayesian solution asymptotically. Moreover, this near-Gaussian form of the IM output offers some valuable insights and simplifications relevant to computing the inner probabilistic approximation; see Section 5. The first order of business is to define Gaussian possibility.
Let $g_{m,v}$ denote the $D$-dimensional Gaussian probability density function, parametrized by a mean vector $m\in\mathbb{R}^{D}$ and a covariance matrix $v\in\mathbb{R}_{+}^{D\times D}$. Define the corresponding Gaussian possibility measure as the outer possibilistic approximation of ${\sf N}_{D}(m,v)$, the probability-to-possibility transform applied to the aforementioned Gaussian distribution. The corresponding possibility contour is
$\gamma_{m,v}(y)=\mathsf{P}\{g_{m,v}(Y)\leq g_{m,v}(y)\},\quad Y\sim{\sf N}_{D}(m,v),$
which can easily be rewritten as
@@
where $F_{D}$ is the ${\sf ChiSq}(D)$ distribution function. Write $\mathsf{\Gamma}_{m,v}$ for the corresponding possibility measure obtained by maximizing the contour as described above. Other authors have used a different form of Gaussian possibility (e.g., Denœux, 2023b ; Denœux, 2023a ), but I find the one above appealing and convenient because of the following duality between ${\sf N}_{D}(m,v)$ and $\mathsf{\Gamma}_{m,v}$: the former is the inner probabilistic approximation of the latter, and the latter is the outer possibilistic approximation of the former.
For some intuition about how this is related to the possibilistic IMs, consider a simple Gaussian location problem, $X\sim\mathsf{P}_{\Theta}={\sf N}_{D}(\Theta,v)$, with known covariance matrix $v$. Then the relative likelihood is $R(x,\theta)=\exp\{-\frac{1}{2}(x-\theta)^{\top}v^{-1}(x-\theta)\}$ and the IM’s contour is
$\pi_{x}(\theta)$ $=\mathsf{P}_{\theta}\{R(X,\theta)\leq R(x,\theta)\}$
$=1-F_{D}\bigl((x-\theta)^{\top}v^{-1}(x-\theta)\bigr),\quad\theta\in\mathbb{R}^{D}.$
@@
$\pi_{x^{n}}(\theta)\approx 1-F_{D}\bigl((\hat{\theta}_{x^{n}}-\theta)^{\top}J_{x^{n}}(\hat{\theta}_{x^{n}}-\theta)\bigr),$
where $\hat{\theta}_{x^{n}}$ is the maximum likelihood estimator and $J_{x^{n}}$ is the observed Fisher information matrix. The reader will notice that the right-hand side above is exactly the Gaussian possibility contour with mean vector $\hat{\theta}_{x^{n}}$ and covariance matrix $J_{x^{n}}^{-1}$, hence the connection between the possibilistic IM and the Gaussian possibility is very natural and holds quite generally, at least approximately. Making this connection rigorous requires care, both in formulating the proper regularity conditions and in suitably bounding all the error terms. Along these lines, under regularity conditions comparable to those advanced in Le Cam, (1970), Martin and Williams, (2025) proved the following Bernstein–von Mises theorem for the possibilistic IM.
@@
Under the regularity conditions presented in Appendix D, as $n\to\infty$, the possibilistic IM’s contour $\pi_{X^{n}}$ satisfies
$\sup_{\theta\in\mathcal{T}}\bigl|\pi_{X^{n}}(\theta)-\gamma_{X^{n}}(\theta)\bigr|\to 0\quad\text{in $\mathsf{P}_{\Theta}$-probability},$
where $\mathcal{T}$ is a fixed but arbitrary compact subset of $\mathbb{T}$,
$\gamma_{X^{n}}(\theta):=\gamma_{\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1}}(\theta)$
is the centered and scaled Gaussian possibility contour, $I_{\Theta}$ the Fisher information matrix, and $\Delta_{\Theta}(X^{n})$ a sequence that converges in distribution to ${\sf N}_{D}(0,I_{\Theta}^{-1})$. In addition, if the likelihood is twice differentiable, then, with $\mathsf{P}_{\Theta}$-probability converging to 1 as $n\to\infty$,
$\sup_{u\in\mathcal{U}}\bigl|\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}u)-\gamma(u)\bigr|\to 0,$
@@
Among other things, Theorem 4 implies consistency: $\mathsf{\Pi}_{X^{n}}(H)\to 0$ in $\mathsf{P}_{\Theta}$-probability for any $H$ with $\Theta\not\in H$. Moreover, since the covariance matrix in the Gaussian limit agrees with the Cramér–Rao lower bound, Theorem 4 implies that the possibilistic IM is asymptotically efficient and, hence, there’s no loss of efficiency associated with the IM’s exact validity (and imprecision). Martin and Williams, (2025) give numerical illustrations that highlight the accuracy of the Gaussian approximation.
A direct consequence of Theorem 4 is that the IM’s inner probabilistic approximation converges to a Gaussian probability distribution just like in the classical, Bayesian Bernstein–von Mises theorem. The proof is in Appendix D of the supplementary material.
@@
Under the conditions of Theorem 4, the proposed inner probabilistic approximation $\mathsf{Q}_{X^{n}}^{\star}$ of the likelihood-based possibilistic IM satisfies a probabilistic Bernstein–von Mises theorem, i.e.,
$d_{\text{\sc tv}}(\mathsf{Q}_{X^{n}}^{\star},\mathsf{G}_{X^{n}}):=\sup_{H}\bigl|\mathsf{Q}_{X^{n}}^{\star}(H)-\mathsf{G}_{X^{n}}(H)\bigr|\to 0,$
where the convergence is in $\mathsf{P}_{\Theta}$-probability as $n\to\infty$, where
$\mathsf{G}_{X^{n}}={\sf N}_{D}(\Theta+n^{-1/2}\Delta_{\Theta}(X^{n}),(nI_{\Theta})^{-1})$
is the approximating Gaussian distribution, and the supremum is over all Borel measurable subsets $H$ of $\mathbb{T}$.
I’ll give a quick proof sketch for a scalar $\Theta$. As in Section 3.3, using $w=\frac{1}{2}$, the distribution function of $\mathsf{Q}_{x^{n}}^{\star}$ is
$\mathsf{Q}_{x^{n}}^{\star}\bigl((-\infty,\theta]\bigr)$ $=\begin{cases}\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta\leq\hat{\theta}_{x^{n}}$}\\ 1-\tfrac{1}{2}\,\pi_{x^{n}}(\theta)&\text{if $\theta>\hat{\theta}_{x^{n}}$}.\end{cases}$
Switching to local coordinates, i.e., $\theta=\hat{\theta}_{x^{n}}+J_{x^{n}}^{-1/2}z$ for $z$ fixed, the case $\theta\leq\hat{\theta}_{x^{n}}$ corresponds to $z\leq 0$ and Theorem 4 implies that, with $Z\sim{\sf N}(0,1)$,
@@
$=\tfrac{1}{2}\,\mathsf{P}(Z^{2}\geq z^{2})$
@@
Similarly, if $z>0$, then $1-\tfrac{1}{2}\,\pi_{X^{n}}(\hat{\theta}_{X^{n}}+J_{X^{n}}^{-1/2}z)\to\mathsf{P}(Z\leq z)$. Therefore, with probability tending to 1, the distribution of $J_{X^{n}}^{1/2}(\Theta-\hat{\theta}_{X^{n}})$, under $\Theta\sim\mathsf{Q}_{X^{n}}^{\star}$, is converging to a standard normal, just like in the classical Bernstein–von Mises theorem.
A point that follows immediately from the above sketch is that assigning probability $\frac{1}{2}$ to the boundary points in the inner probabilistic approximation is crucial; of course, any choice that assigns probability to the boundary points that converges to $\frac{1}{2}$ works fine too. More generally, since $\mathsf{G}_{X^{n}}$ is an inner probabilistic approximation to the possibility measure with contour $\gamma_{X^{n}}$, and corresponds to taking a uniform kernel on the level sets of $\gamma_{X^{n}}$, it’s necessary that the choice of kernel defining $\mathsf{Q}_{X^{n}}^{\star}$ at least be asymptotically uniform. This observation adds further support for my recommendation above to adopt the uniform kernel in the inner probabilistic approximation.